# OralBBNet: Spatially Guided Dental Segmentation of Panoramic X-Rays with Bounding Box Priors

### v2

@@ Abstract
Teeth segmentation and recognition play a vital role in a variety of dental applications and diagnostic procedures. The integration of deep learning models has facilitated the development of precise and automated segmentation methods. Although prior research has explored teeth segmentation, not many methods have successfully performed tooth segmentation and detection simultaneously. This study presents UFBA-425, a dental dataset derived from the UFBA-UESC dataset, featuring bounding box and polygon annotations for $425$ panoramic dental X-rays. Additionally, this work introduces OralBBNet, an architecture featuring distinct segmentation and detection heads as U-Net and YOLOv8, respectively. OralBBNet is designed to improve the accuracy and robustness of tooth classification and segmentation on panoramic X-rays by leveraging the complementary strengths of U-Net and YOLOv8. Our approach achieved a $1-3\%$ improvement in mean average precision (mAP) for teeth detection compared to existing techniques and a $15-20\%$ improvement in the dice score for teeth segmentation over U-Net over various tooth categories and $2-4\%$ improvement in the dice score when compared with other segmentation architectures. The results of this study establish a foundation for the wider implementation of object detection models in dental diagnostics.
@@ 1.1 Motivation
Given these challenges, automated deep learning-based teeth segmentation and classification have gained increasing attention due to their success in handling complex computer vision tasks. However, existing approaches still face difficulties in precisely recognizing and localizing each tooth due to the aforementioned issues. Thus, improving the robustness and accuracy of automated segmentation methods remains a critical research challenge.
@@ 1.2 Contributions
This work presents OralBBNet, an enhanced method for classifying teeth and performing instance segmentation, which incorporates spatial prior knowledge into a U-Net [6] framework. By employing a one-stage detection head, YOLO [7], to capture spatial features, we improve both efficiency and accuracy of the segmentation, moving away from traditional two-stage object detection models like Mask R-CNN. The major contributions of our work are:
@@
    OralBBNet, a new architecture, was developed to perform both teeth numbering and segmentation with improved spatial prior knowledge.
@@ 3.3 Model Architecture
3.3.1 Detection Head: YOLO Architecture
@@
3.3.2 Segmentation Head: U-Net Architecture
@@ 3.3.3 OralBBNet Architecture
OralBBNet builds upon the U-Net framework by incorporating bounding box priors produced by the detection head into the learning process, enhancing segmentation precision through explicit spatial guidance. It’s key innovation is the BB-Convolution layers, which refine feature maps using bounding-box-based supervision. While the encoder maintains U-Net’s hierarchical convolutional design, it integrates bounding-box-driven refinements at multiple levels. These bounding boxes act as spatial anchors, directing feature extraction to reduce errors in segmenting adjacent or overlapping structures.
@@ 3.4 Training Pipeline and Hyperparameters Setup
We propose a pipeline featuring a two-stage procedure that involves training the YOLOv8 and OralBBNet architectures independently, as illustrated in Figure  5. In the first stage, the YOLOv8 detection head is trained to identify teeth, determine their positions, and extract bounding boxes. In the subsequent stage, the trained YOLOv8 model provides spatial prior information of bounding boxes, serving as a detection head while training OralBBNet for instance segmentation, ultimately producing segmentation masks as output. Additionally, we ensured that the weights of YOLOv8 remained unchanged during the second stage of training for OralBBNet to avoid collapse of weights in YOLOv8.
@@
3.4.1 Stage 1: Training of Detection Head
@@
3.4.2 Stage 2: Training of Segmentation Head
@@ 4.1 Evaluation Metrics
We have used several metrics to evaluate the quality of the proposed method. We calculate accuracy, recall, precision, and mAP:
@@
4.2 Evaluation of Detection Head: YOLOv8
@@ 4.2.2 Qualitative Analysis
4.3 Evaluation of Segmentation Head: OralBBNet
@@ 4.3.2 Qualitative Analysis
OralBBNet segmentation method notably outperforms the standard U-Net segmentation regarding quality. In Figure 10, the top row highlights the superior pixel-level classification of teeth achieved by OralBBNet, especially in complex tooth structures like molars and premolars, where U-Net encounters challenges. Nonetheless, the bottom row depicts the scenarios where bounding box predictions are degraded, OralBBNet’s performance aligns with U-Net due to inadequate spatial prior knowledge. Another observation was that OralBBNet successfully predicted most of the pixel densities even in category-5 and category-6 X-ray images but it struggled to accurately delineate the boundary pixels, which limits the dentists to extract precise spatial information from the segmentation masks Moreover, the dependence of OralBBNet’s performance on bounding box predictions emphasizes the importance of comprehensive data annotation and robust performance of detection head.
@@ 5 Comparison Studies
5.1 Comparison of Detection Head with other Tooth Detection Models
@@
5.2 Comparison of Segmentation Head with other Segmentation Models

### v3

@@ Abstract
Teeth segmentation and recognition play a vital role in a variety of dental applications and diagnostic procedures. The integration of deep learning models has facilitated the development of precise and automated segmentation methods. Although prior research has explored teeth segmentation, not many methods have successfully performed tooth segmentation and detection simultaneously. This study presents UFBA-425, a dental dataset derived from the UFBA-UESC dataset, featuring bounding box and polygon annotations for $425$ panoramic dental X-rays. In addition, this paper presents the OralBBNet architecture, which is based on the best segmentation and detection qualities of architectures such as U-Net and YOLOv8, respectively. OralBBNet is designed to improve the accuracy and robustness of tooth classification and segmentation on panoramic X-rays by leveraging the complementary strengths of U-Net and YOLOv8. Our approach achieved a $1-3\%$ improvement in mean average precision (mAP) for tooth detection compared to existing techniques and a $15-20\%$ improvement in the dice score for teeth segmentation over state-of-the-art (SOTA) solutions for various tooth categories and $2-4\%$ improvement in the dice score compared to other SOTA segmentation architectures. The results of this study establish a foundation for the wider implementation of object detection models in dental diagnostics.
@@ 1.1 Motivation
Given these challenges, automated deep learning-based teeth segmentation and classification have gained increasing attention due to their success in handling complex computer vision tasks. However, existing approaches still face difficulties in precisely recognizing and localizing each tooth due to the aforementioned issues. Thus, improving the robustness and accuracy of automated segmentation models remains a critical research challenge.
@@ 1.2 Contributions
This work presents OralBBNet, an enhanced model for classifying teeth and performing instance segmentation, which incorporates spatial prior knowledge into a U-Net [6] framework. By employing a one-stage detection module, YOLO [7], to capture spatial features, we improve both efficiency and accuracy of the segmentation, moving away from traditional two-stage object detection models like Mask R-CNN. The major contributions of our work are:
@@
    OralBBNet, a new segmentation architecture, was developed to perform both teeth numbering and segmentation with improved spatial prior knowledge.
@@ 3.3 Model Architecture
3.3.1 Detection Module: YOLO-based Architecture
@@
3.3.2 Segmentation Module: U-Net-based Architecture
@@ 3.3.3 OralBBNet Architecture
OralBBNet builds on the U-Net framework by incorporating bounding box priors produced by the detection module into the learning process, enhancing segmentation precision through explicit spatial guidance. It’s key innovation is the BB-Convolution layers, which refine feature maps using bounding-box-based supervision. While the encoder maintains U-Net’s hierarchical convolutional design, it integrates bounding-box-driven refinements at multiple levels. These bounding boxes act as spatial anchors, directing feature extraction to reduce errors in segmenting adjacent or overlapping structures.
@@ 3.4 Training Pipeline and Hyperparameters Setup
We propose a pipeline featuring a two-stage procedure that involves training the YOLOv8 and OralBBNet architectures independently, as illustrated in Figure  5. In the first stage, the YOLOv8 detection module is trained to identify teeth, determine their positions, and extract bounding boxes. In the subsequent stage, the trained YOLOv8 model provides spatial prior information of bounding boxes, serving as a detection module while training OralBBNet for instance segmentation, ultimately producing segmentation masks as output. Furthermore, we ensured that the weights of YOLOv8 remained unchanged during the second stage of OralBBNet training to avoid collapse of weights in YOLOv8.
@@
3.4.1 Stage 1: Training of Detection Module
@@
3.4.2 Stage 2: Training of Segmentation Module
@@ 4.1 Evaluation Metrics
We have used several metrics to estimate the quality of the proposed model. We calculate the following metrics:
@@
4.2 Evaluation of Detection Module: YOLOv8
@@ 4.2.2 Qualitative Analysis
4.3 Evaluation of Segmentation Module: OralBBNet
@@ 4.3.2 Qualitative Analysis
OralBBNet segmentation model notably outperforms the standard U-Net-based segmentation models from the point of quality. In Figure 10, the top row highlights the superior pixel-level classification of teeth achieved by OralBBNet, especially in complex tooth structures like molars and premolars, where U-Net encounters challenges. Nonetheless, the bottom row depicts the scenarios where bounding box predictions are degraded, OralBBNet’s performance aligns with U-Net due to inadequate spatial prior knowledge. Another observation was that OralBBNet successfully predicted most of the pixel densities even in category-5 and category-6 X-ray images but it struggled to accurately delineate the boundary pixels, which limits the dentists to extract precise spatial information from the segmentation masks. Moreover, the dependence of OralBBNet’s performance on bounding box predictions emphasizes the importance of comprehensive data annotation and robust performance of detection module.
@@ 5 Comparison Studies
5.1 Comparison of Detection Module with other Tooth Detection Models
@@
5.2 Comparison of Segmentation Module with other Segmentation Models